When 15.0 mL of a 4.80×10-4 M iron(II) bromide solution is combined with 15.0 mL of a 3.90×10-4 M potassium carbonate solution does a precipitate form? (yes or no) For these conditions the Reaction Quotient, Q, is equal to
Answer #1

Answer

Yes, precipitate will form

Q = 4.68 ×10-8

Explanation

After mixing the solution

[Fe2+] = 4.80 ×10-4M/ 2 = 2.40×10-4M

[CO32-] = 3.90 ×10-4M/2 = 1.95 ×10-4M

Q= 2.40 × 10-4M × 1.95 ×10-4M = 4.68 ×10-8

Ksp of FeCO3 = 3.13×10-11

Q > Ksp ,

Therefore

precipitate will form